# seq-JEPA: Autoregressive Predictive Learning of Invariant-Equivariant World Models

### v1

@@ Abstract
Current self-supervised algorithms mostly rely on transformations such as data augmentation and masking to learn visual representations. This is achieved by inducing invariance or equivariance with respect to these transformations after encoding two views of an image. This dominant two-view paradigm can limit the flexibility of learned representations for downstream adaptation by creating performance trade-offs between invariance-related tasks such as image classification and more fine-grained equivariance-related tasks. In this work, we introduce seq-JEPA, a world modeling paradigm based on joint-embedding predictive architecture that leverages architectural inductive biases to resolve this trade-off. Without requiring an additional equivariance predictor or loss term, seq-JEPA simultaneously learns two architecturally segregated representations: one equivariant to the specified transformations and another invariant to them and suited for tasks such as classification. To do so, our model processes a short sequence of different views (observations) of an input image. Each encoded view is concatenated with embeddings corresponding to the relative transformation (action) producing the next observation in the sequence. A transformer encoder outputs an aggregate representation of this sequence, which is subsequently conditioned on the action leading to the next observation to predict its representation. Empirically, seq-JEPA achieves strong performance on equivariant benchmarks and image classification without sacrificing one for the other. Additionally, our framework excels at tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
Self-supervised learning, world models, computer vision
Figure 1: Top-1 linear classification (invariance) versus rotation prediction (equivariance) performance on 3DIEBench. Our world model, seqJEPA, learns representations for both invariance- and equivariance-related tasks at the same time. Subscripts indicate training and inference sequence lengths used in seq-JEPA.
@@ 1 Introduction
Self-supervised learning (SSL) in latent space has made significant progress in visual representation learning in recent years, closing the gap with supervised learning in many downstream tasks (van den Oord et al., 2019; Misra & van der Maaten, 2020; He et al., 2020; Chen et al., 2020; Grill et al., 2020; Chen & He, 2021; Caron et al., 2020, 2021; Zbontar et al., 2021; Bardes et al., 2022; Baevski et al., 2022; Assran et al., 2023). Most SSL methods learn their representations by comparing two transformed views of a sample image and inducing invariance with respect to these transformations (Chen et al., 2020; He et al., 2020; Dwibedi et al., 2021; HaoChen et al., 2021; Yeh et al., 2022; Caron et al., 2020, 2021; Ermolov et al., 2021; Assran et al., 2022; Zbontar et al., 2021; Bardes et al., 2022). Another group of methods employ techniques to preserve some information about the transformations and learn equivariant representations (Lee et al., 2021; Xiao et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024).
@@
Equivariance is a required representational property for many fine-grained downstream tasks. For example, given representations that are invariant to color, it is not possible to distinguish between certain species of flowers or birds (Lee et al., 2021; Xiao et al., 2021). Additionally, recent work has shown that equivariant representations are better aligned with neural responses in primate visual cortex and could be important for building more accurate models thereof (Yerxa et al., 2024). Although some of the equivariant SSL methods have empirically shown a slight performance improvement in downstream tasks traditionally associated with invariance (e.g., image classification) (Devillers & Lefort, 2022; Park et al., 2022; Gupta et al., 2023), recent studies have revealed a persistent trade-off between invariance and equivariance in self-supervised representation learning (Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024). The existence of such a trade-off has been recently supported by theoretical arguments (Wang et al., 2024), highlighting the need for innovative architectural and objective design to mitigate it.
@@
Humans and other animals rely on a sequence of actions and consequent observations (views) for developing appropriate visual representation during novel object learning  (Harman et al., 1999; Vuilleumier et al., 2002). For example, they recognize a 3-D object by changing their viewpoint and examining different sides of the object (Tarr et al., 1998). Inspired by this, we introduce seq-JEPA, a world modeling paradigm based on joint-embedding predictive architecture (LeCun, 2022; Dawid & LeCun, 2024) that leverages architectural inductive biases to enable such sequential processing in SSL and resolve the invariance-equivariance trade-off. Our model simultaneously learns two architecturally distinct representations: one equivariant to a specified set of transformations and another suited for invariance-related tasks, such as image classification.
@@
Specifically, our framework (Figure 2) processes a short sequence of different views (observations) of an input image. Each encoded view is concatenated with embeddings corresponding to the relative transformation (action) that produces the next observation in the sequence. These concatenated representations are then fed to a transformer encoder, resembling working memory, that outputs an aggregate representation, which is subsequently conditioned on the action leading to the next observation and used to predict its representation.
@@
Our results show that the encoded views in seq-JEPA become transformation/action-equivariant. Through ablation experiments, we show that action conditioning plays a key role in promoting equivariant representation learning in the encoder network. In contrast, the aggregate representation of views, produced at the output of the sequence transformer, becomes largely transformation/action-invariant. This architectural disentanglement of invariant and equivariant representations in seq-JEPA is key to the competitive performance of our model compared to both invariant and equivariant SSL methods (Figure 1). We empirically demonstrate this advantage across multiple tasks and benchmarks. It is important to emphasize that, unlike many previous approaches in equivariant SSL (Lee et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023; Gupta et al., 2024; Yerxa et al., 2024), neither our objective function nor our architecture explicitly encourages the learning of invariant-equivariant representations. Instead, the learning of invariance and equivariance, along with their architectural separation, emerges as an implicit consequence of action-conditioned sequential predictive learning and the model architecture.
@@
Furthermore, we explore other benefits of moving away from the dominant two-view paradigm in SSL and demonstrate that seq-JEPA is capable of solving tasks that inherently require aggregating a sequence of observations. In one scenario, inspired by embodied vision in primates, our model learns image representations without hand-crafted augmentations or masking, solely via predictive learning across simulated saccades (eye movements). Finally, seq-JEPA is also capable of path integration over a sequence of actions such as eye movements in the previously described scenario, or 3-D object rotations in 3DIEBench (Garrido et al., 2023). A summary of our contributions are as follows:
  • We propose seq-JEPA, a self-supervised world model that leverages a sequence of action-observation pairs to learn separate representations for equivariance- and invariance-demanding downstream tasks without requiring an additional equivariance predictor or loss term.
  • We empirically show the effectiveness of our world modeling paradigm in comparison with invariant and equivariant SSL methods on both invariance- and equivariance-related benchmarks without sacrificing one for the other.
@@
  • We demonstrate that seq-JEPA is capable of solving tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
@@ 2.1 Invariant and equivariant representations
Before describing the architecture and training procedure of seq-JEPA, we provide the nomenclature of invariant and equivariant representation learning in SSL (Dangovski et al., 2022; Devillers & Lefort, 2022). Consider $\mathcal{T}$ as a distribution of possible data transformations that is parametrized by a vector $t$. These transformations can be any type of hand-crafted data augmentation, such as geometric transformations or masking, that transform an original data point $x$ to different views. For example, we denote $x_{1}$ and $x_{2}$ as views generated from $x$ via two transformations $t_{1}$ and $t_{2}$ sampled from $\mathcal{T}$. Additionally, we denote $a$ as a transformation that transforms $x_{1}$ to $x_{2}$. Here, we distinguish between $t$ and $a$, and call $t$ a transformation and $a$ an action, i.e., a relative transformation that transforms $x_{1}$ to $x_{2}$. Furthermore, we denote $f$ as an encoder function that projects the data space into a latent space. Given a transformation $t$, the encoder and its representations are said to be equivariant to $t$ if
$\forall{t\in\mathcal{T}},~{}\exists{\,u_{t}}\;\mathrm{s.t.}\qquad f(t(x))=u_{t}(f(x)).$ (1)
We can also define equivariance given the action $a$ and two views $x_{1}$ and $x_{2}$ as
$\forall{a\in\mathcal{T}},~{}\exists{\,u_{a}}\;\mathrm{s.t.}\qquad f(x_{2})=u_{a}(f(x_{1})),$ (2)
where $u_{t}$ and $u_{a}$ are functions that produce displacements in latent space corresponding to their associated transformation/action. As a special case of equivariance, the encoder and its representations are said to be invariant to $t$ if $u_{t}$ and $u_{a}$ are identity functions, i.e., $f(t(x))=f(x)$ or $f(x_{2})=f(x_{1})$.
@@ 2.2 Architecture
The overall architecture of seq-JEPA is depicted in Figure 2. Denote $\{x_{i}\}_{i=1}^{M+1}$ as a sequence of views generated from a data sample $x$ using transformations $\{t_{i}\}_{i=1}^{M+1}$. The corresponding actions (relative transformations), whose embeddings are learned via a linear layer are denoted as $\{a_{i}\}_{i=1}^{M}$. A backbone encoder, $f_{\theta}$ encodes the first $M$ views and produces representations $\{z_{i}\}_{i=1}^{M}$. These representations, except for $z_{M}$, are concatenated with their corresponding action embeddings and fed to a sequence model, e.g., a transformer encoder, $g_{\phi}$ (no MLP projector is used after the encoder). This transformer module is tasked with aggregating the sequence of representations of action-observation pairs to produce an aggregate representation. We use a learnable token which we call [AGG] to learn the aggregate representation (this token resembles a [CLS] token in vision transformers (Dosovitskiy et al., 2020)). The output corresponding to [AGG] is therefore:
$z_{AGG}=g_{\phi}((z_{1},a_{1}),(z_{2},a_{2})...,(z_{M-1},a_{M-1}),z_{M})$ (3)
The representation $z_{AGG}$ is subsequently concatenated with the last action embedding $a_{M}$ which corresponds to the relative transformation between $x_{M}$ and $x_{M+1}$, and is fed to an MLP predictor module $h_{\psi}$ to predict the representation of $x_{M+1}$, i.e., $\hat{z}_{M+1}=h_{\psi}(z_{AGG},a_{M})$. The last view $x_{M+1}$, is itself encoded using a target encoder whose weights are an exponential moving average (EMA) of the encoder $f_{\theta}$ used to encode the sequence of observations. The representation $z_{M+1}$ is passed through an stop-gradient (sg) to prevent representational collapse. Our objective is then to maximize the cosine similarity between $\hat{z}_{M+1}$ and $z_{M+1}$, which we achieve by minimizing the following prediction loss,
$\mathcal{L}_{seq-JEPA}=1-\frac{{\hat{z}_{M+1}}}{\left\lVert{{\hat{z}_{M+1}}}\right\rVert_{2}}\cdot\frac{{\texttt{sg}(z_{M+1}})}{\left\lVert{{\texttt{sg}(z_{M+1}})}\right\rVert_{2}}.$ (4)
No additional equivariance predictor or loss function is used when training the framework.
Figure 2: Overview of the proposed architecture; seq-JEPA is a world modeling paradigm that leverages a sequence of action-observation pairs to learn separate representations for equivariance- and invariance-related tasks.
@@ 2.3 Action and observation sets
In order to evaluate our model and measure its ability to generalize to different types of transformations, we consider three distinct sets of transformation and their corresponding action-observation pairs (Figure 3a). Descriptions of each transformation set are provided below, with additional details available in Section A.1.
@@
3D Invariant Equivariant Benchmark (3DIEBench). We use the 3DIEBench dataset (Garrido et al., 2023) as our first evaluation benchmark. This dataset is specifically designed for measuring equivariance and provides 3D object renderings with different object rotations and floor hue and lighting angles and their corresponding parameters. We specifically study equivariance to the 3D rotation group ($SO(3)$) and further to floor and light hue in this benchmark.
@@
Hand-crafted data augmentations. In this setting, observations are transformed image views generated by hand-crafted augmentations that are typically used in the two-view paradigm of SSL. The actions are relative augmentation parameters between two views. We measure equivariance to crop, color jitter, and blur transformations. As benchmarks for this setting, we use CIFAR100 and Tiny ImageNet datasets. To generate augmented views, we follow EquiMod’s augmentation strategy (Devillers & Lefort, 2022). For details on augmentations and the corresponding parameterization, see Section A.1.
Predictive learning across saccades. Going beyond transformations such as augmentations and rotation, we explore other ways of leveraging sequential observations to learn visual representations. Specifically, we designed a procedure to learn image representations without using any hand-crafted augmentations with a similar flavor to I-JEPA (Assran et al., 2023) but without dependence on masking. Our approach is enabled by seq-JEPA’s architecture and its ability to produce an aggregate representations from a sequence of observations. In this setting, given a full-size high-resolution image, we extract a number of smaller patches and use them as a sequence of partial observations. In our experiments with the STL-10 dataset, we use 32 by 32 patches, i.e. one-ninth of the original image.
The action in this setup corresponds to the relative position between the center of two subsequent patches, i.e. a simulated act of rapid eye movements (also known as saccades) from one point of the image to another. Although it is possible to use random positions as patch centers (fixations) for the sequence, we show that adopting two techniques inspired by embodied vision in primates improves the quality of the aggregate representation and its downstream classification performance (Figure 3b). The first technique utilizes image saliencies (Itti et al., 1998; Li, 2002; Zhaoping, 2014) to sample fixations. To generate saliency maps, we use a pre-trained saliency prediction model, DeepGaze IIE (Linardos et al., 2021). Given a dataset, we pre-extract and store the saliencies without any additional overhead during training. These saliency maps can be interpreted as how likely humans are to center their gaze (fixate) on a given pixel of the image. Therefore, sampling fixations from the probability distribution corresponding to saliency maps provides more informative patches, yet similar to uniform sampling, is prone to a high degree of redundancy and overlap between patches. To remedy this, i.e. minimize overlap among patches and avoid future fixation that revisit previously attended areas, we use another technique inspired by biological vision known as inhibition of return (IoR) (Posner et al., 1985). To implement IoR, after sampling a fixation from saliency maps, we zero out the sampling probabilities of an area with a diameter equal to the patch width surrounding the previous fixations before sampling a subsequent fixation. In this scenario, we are interested to see whether encoder representations become sensitive to fixation positions (spatial awareness), and therefore equivariance is measured with respect to fixation coordinates.
Figure 3: a. Types of transformations and observations used for training; seq-JEPA works with a variety of action-observation pairs such as 3D object rotations, hand-crafted augmentations, and simulated saccades. b. For predictive learning across saccades, image saliencies and inhibition of return help create more informative and less redundant patch sequences.
3 Related Work
Joint-embedding predictive architectures. In JEPAs (LeCun, 2022; Dawid & LeCun, 2024), a predictor asymmetry is introduced in one of the two branches that traditionally encode transformed views in SSL. The representation learning task is then framed as a world modeling objective (Ha & Schmidhuber, 2018; Hafner et al., 2019), i.e., the predictor is tasked with predicting the consequence of a data transformation (action) applied to the first view that results in the second view. The predictor is conditioned on transformation parameters such as positional masks to predict the latent representation of target blocks corresponding to the masks as in I-JEPA (Assran et al., 2023), or additionally on parameters corresponding to hand-crafted image augmentations as in Image World Model (IWM) (Garrido et al., 2024). Recent works have also explored JEPAs for physical understanding (Garrido et al., 2025) and reward-free offline planning (Sobal et al., 2025).
Non-generative world models. Non-generative world models predict the consequence of an action or transformation in latent space. A line of work considers such world models in contrastive SSL by predicting the representation of a transformed view from another via a contrastive objective and conditioned on the past or context (van den Oord et al., 2019; Gupta et al., 2024). The family of JEPAs defer the task of world modeling to a predictor module that is conditioned on the action (Assran et al., 2023; Garrido et al., 2024; Sobal et al., 2025). Non-generative world models have also garnered attention in model-based reinforcement learning (RL) and planning to improve sample efficiency (Schwarzer et al., 2021), to learn environment dynamics (Sobal et al., 2025), or to create an intrinsic novelty-based reward to improve exploration in self-predictive world models (Khetarpal et al., 2025; Ni et al., 2024; Tang et al., 2023). A prominent example of such world models is BYOL-Explore (Guo et al., 2022), in which a recurrent cell follows the encoder and is conditioned on the current action in a planning task, followed by a predictor to capture the environment’s transition dynamics.
Equivariant representation learning. Many equivariant SSL methods extend invariant methods by introducing an additional equivariance objective. This objective can be direct prediction of parameters corresponding to augmented views (Lee et al., 2021; Scherr et al., 2022; Gidaris et al., 2018; Gupta et al., 2024), or additional transformations such as rotation (Dangovski et al., 2022). EquiMod (Devillers & Lefort, 2022) and SIE (Garrido et al., 2023) predict the effect of a transformation in latent space in addition to their invariant objective. Xiao et al. (2021) use contrastive learning to train multiple projection heads for equivariance to each augmentation, where the views corresponding to the desired augmentation are considered as negative samples. In SEN (Park et al., 2022), similar to EquiMod, a predictor is used to predict the effect of transformations in latent space, yet without any invariance objective. ContextSSL (Gupta et al., 2024) learns representations that are equivariant only to the recent transformations in the current context by conditioning representations on both the current action and the past via a memory module, and further using a transformation prediction objective to avoid collapse to invariance. Another line of work including Shakerinava et al. (2022), CARE (Gupta et al., 2023), and Yerxa et al. (2024) do not assume access to transformation parameters yet require two pairs of views to which the same transformation is applied, and an equivariance loss function is minimized to induce equivariance in representations. Generative SSL methods based on masked image modeling (He et al., 2022; Bao et al., 2021; Xie et al., 2022) can become equivariant to position due to the use of positional embeddings. Finally, action-conditioned JEPAs condition the predictor on positional mask tokens (Assran et al., 2023) and augmentation parameters (Garrido et al., 2024) to induce equivariance in encoder representations without any additional equivariance objective.
Relationship to existing literature. Our framework belongs to the family of JEPAs and is a non-generative world model. We show that via architectural inductive biases, JEPAs can be leveraged to resolve the invariance-equivariance trade-off in SSL and enable processing of sequential action-observation pairs that is required to solve specific perception tasks. Unlike the majority of equivariant SSL methods discussed above, seq-JEPA does not use an equivariance predictor or loss function, nor does it require two pairs of views to which the same transformation is applied. We further highlight the differences between our framework and a recent equivariant method, ContextSSL. Specifically, seq-JEPA can be distinguished from ContextSSL on two aspects. First, unlike seq-JEPA, ContextSSL operates within the two-view SSL paradigm but uses a decoder-only transformer instead of the usual MLP projector so that its output embeddings are conditioned on the past. Second, while ContextSSL’s main goal is adapting equivariance to different transformations by alternating transformations in training data, our goal is to learn separate invariant and equivariant representations with respect to one or more transformations specified, and solve downstream tasks that require aggregating a sequence of action-observation pairs.

### v2

@@ Abstract
Current self-supervised algorithms commonly rely on transformations such as data augmentation and masking to learn visual representations. This is achieved by enforcing invariance or equivariance with respect to these transformations after encoding two views of an image. This dominant two-view paradigm often limits the flexibility of learned representations for downstream adaptation by creating performance trade-offs between high-level invariance-demanding tasks such as image classification and more fine-grained equivariance-related tasks. In this work, we proposes seq-JEPA, a world modeling framework that introduces architectural inductive biases into joint-embedding predictive architectures to resolve this trade-off. Without relying on dual equivariance predictors or loss terms, seq-JEPA simultaneously learns two architecturally segregated representations: one equivariant to specified transformations and another invariant to them. To do so, our model processes short sequences of different views (observations) of inputs. Each encoded view is concatenated with an embedding of the relative transformation (action) that produces the next observation in the sequence. These view-action pairs are passed through a transformer encoder that outputs an aggregate representation. A predictor head then conditions this aggregate representation on the upcoming action to predict the representation of the next observation. Empirically, seq-JEPA demonstrates strong performance on both equivariant and invariant benchmarks without sacrificing one for the other. Furthermore, it excels at tasks that inherently require aggregating a sequence of observations, such as path integration across actions and predictive learning across eye movements.
@@ 1 Introduction
Self-supervised learning (SSL) in latent space has made significant progress in visual representation learning, closing the gap with supervised methods across many tasks. Most SSL methods rely on comparing two transformed views of an image and enforcing invariance to the transformations (Misra and van der Maaten, 2020; Chen et al., 2020; He et al., 2020; Dwibedi et al., 2021; HaoChen et al., 2021; Yeh et al., 2022; Caron et al., 2020, 2021; Ermolov et al., 2021; Assran et al., 2022; Zbontar et al., 2021; Bardes et al., 2022). Another group of methods employ techniques to preserve transformation-specific information, thereby learning equivariant representations (Lee et al., 2021; Xiao et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023, 2024; Gupta et al., 2023, 2024; Yerxa et al., 2024).
@@
Equivariance is a crucial representational property for downstream tasks that require fine-grained distinctions. For example, given representations that are invariant to color, it is not possible to distinguish between certain species of flowers or birds (Lee et al., 2021; Xiao et al., 2021). Moreover, recent work has shown that equivariant representations are better aligned with neural responses in primate visual cortex and could be important for building more accurate models thereof (Yerxa et al., 2024). While some equivariant SSL approaches have reported minor gains on tasks typically associated with invariance (e.g., classification) (Devillers and Lefort, 2022; Park et al., 2022; Gupta et al., 2023), a growing body of work highlights a fundamental trade-off between learning invariance and equivariance, i.e., models that capture equivariance-related style latents do not fare well in classification and vice versa (Garrido et al., 2023, 2024; Gupta et al., 2024; Yerxa et al., 2024; Rusak et al., 2025). This trade-off has recently received theoretical support (Wang et al., 2024), underscoring the need for new architectural and objective designs that can reconcile these competing goals.
@@
In contrast to the two-view paradigm in SSL, humans and other animals rely on a sequence of actions and consequent observations (views) for developing appropriate visual representation during novel object learning  (Harman et al., 1999; Vuilleumier et al., 2002). For example, they recognize a 3-D object by changing their viewpoint and examining different sides of the object (Tarr et al., 1998). Inspired by this, we introduce seq-JEPA, a self-supervised world modeling framework that combines joint-embedding predictive architectures (LeCun, 2022) with inductive biases for sequential processing. seq-JEPA simultaneously learns two architecturally distinct representations: one that is equivariant to a specified set of transformations, and another that is suited for invariance-demanding tasks, such as image classification.
@@
Specifically, our framework (Figure 1) processes a short sequence of transformed views (observations) of an image. Each view is encoded and concatenated with an embedding corresponding to the relative transformation (action) that produces the next observation in the sequence. These view-action pairs are passed through a transformer encoder, a form of learned working memory, that outputs an aggregate representation of them. A predictor head then conditions this aggregate representation on the upcoming action to predict the representation of the next observation.
@@
Our results demonstrate that individual encoded views in seq-JEPA become transformation/action-equivariant. Through ablations, we show that action conditioning plays a key role in promoting equivariant representation learning in the encoder network. In contrast, the aggregate representation of views, produced at the output of the transformer, becomes largely transformation/action-invariant. This emergent architectural disentanglement of invariant and equivariant representations is central to seq-JEPA’s competitive performance compared to invariant and equivariant SSL methods on both categories of tasks (Figure 3). Unlike most prior equivariant SSL methods (Lee et al., 2021; Park et al., 2022; Dangovski et al., 2022; Gupta et al., 2023; Garrido et al., 2023; Gupta et al., 2024; Yerxa et al., 2024), our model does not rely on explicitly crafted loss terms or objectives to achieve equivariance, nor is it instructed to learn the decomposition of two representation types. Instead, the dual representation structure arises naturally from the model architecture and action-conditioned predictive learning.
@@
Beyond resolving the invariance-equivariance trade-off, seq-JEPA further benefits from moving away from the two-view paradigm in SSL. We show that it performs well on tasks requiring integration over sequences of observations. In one scenario, inspired by embodied vision in primates, our model learns image representations without augmentations or masking, solely by predicting across simulated eye movements (saccades). In another setting, it performs path integration over sequences of actions—such as eye movements or 3D object rotations in 3DIEBench (Garrido et al., 2023). Our key contributions are as follows:
  • We introduce seq-JEPA, a self-supervised world model that learns architecturally distinct equivariant and invariant representations through sequential prediction over action-observation pairs, without requiring explicit equivariance losses or dual predictors.
  • We empirically validate that seq-JEPA matches or outperforms existing invariant and equivariant SSL methods across tasks requiring either representational property.
@@
  • We demonstrate that seq-JEPA naturally supports tasks that involve sequential integration of observations, such as predictive learning across saccades and path integration over action sequences.
@@ 2.1 Invariant and equivariant representations
Before presenting our architecture and training procedure, we briefly define invariance and equivariance in the context of SSL (Dangovski et al., 2022; Devillers and Lefort, 2022). Let $\mathcal{T}$ denote a distribution over transformations, parameterized by a vector $t$. These transformations—such as augmentations or masking—can be used to generate multiple views from a single image $x$. Let $x_{1}$ and $x_{2}$ be two such views, produced by applying transformations $t_{1}$ and $t_{2}$ sampled from $\mathcal{T}$. Additionally, let $a$ denote the relative transformation that maps $x_{1}$ to $x_{2}$. Additionally, we denote $a$ as a transformation that transforms $x_{1}$ to $x_{2}$. We distinguish between $t$ (an individual transformation) and $a$ (an action), where the latter reflects the change from one view to another. Let $f$ be an encoder that maps inputs to a latent space. We say that $f$ is equivariant to $t$ if:
$\forall{t\in\mathcal{T}},~{}\exists{,u_{t}};\mathrm{s.t.}\qquad f(t(x))=u_{t}(f(x)),$ (1)
where $u_{t}$ is a transformation in latent space corresponding to $t$. Equivariance can similarly be defined in terms of relative transformations (actions):
$\forall{a\in\mathcal{T}},~{}\exists{,u_{a}};\mathrm{s.t.}\qquad f(x_{2})=u_{a}(f(x_{1})).$ (2)
As a special case, if $u_{t}$ and $u_{a}$ are identity functions, then $f$ is invariant to the transformation: $f(t(x))=f(x)$ or $f(x_{2})=f(x_{1})$.
@@ 2.2 Architecture
Figure 1 presents the overall architecture of seq-JEPA. Let $\{x_{i}\}_{i=1}^{M+1}$ be a sequence of views generated from a sample $x$ via transformations $\{t_{i}\}_{i=1}^{M+1}$. The relative transformations (actions) $\{a_{i}\}_{i=1}^{M}$ are defined as $a_{i}\triangleq\Delta t_{i,i+1}$, i.e., the transformation mapping $x_{i}$ to $x_{i+1}$. In our default setting, we use a learnable linear projector to encode these actions.
A backbone encoder, $f$ encodes the first $M$ views, producing representations $\{z_{i}\}_{i=1}^{M}$. Except for $z_{M}$, each $z_{i}$ is concatenated with its corresponding action embedding and passed to a transformer encoder $g$ (no MLP projector is used after the encoder), which aggregates the sequence of action-observation pairs. The transformer uses a learnable [AGG] token (analogous to the [CLS] token in ViTs (Dosovitskiy et al., 2020)) to generate the aggregate representation:
$z_{AGG}=g((z_{1},a_{1}),(z_{2},a_{2})...,(z_{M-1},a_{M-1}),z_{M})$ (3)
This aggregate representation $z_{AGG}$ is then concatenated with the final action embedding $a_{M}$ (corresponding to the transformation from $x_{M}$ to $x_{M+1}$), and passed to an MLP predictor $h$ to predict the representation of $x_{M+1}$:
$\hat{z}_{M+1}=h(z_{AGG},a_{M}).$ (4)
The ground truth $z_{M+1}$ is computed using a target encoder—an exponential moving average (EMA) of $f$. The target representation is passed through a stop-gradient operator (sg) to avoid representational collapse. The training objective is to maximize the cosine similarity between $\hat{z}_{M+1}$ and $z_{M+1}$:
$\mathcal{L}_{\mathrm{seq-JEPA}}=1-\frac{\hat{z}_{M+1}}{\left\lVert{\hat{z}_{M+1}}\right\rVert_{2}}\cdot\frac{\texttt{sg}(z_{M+1})}{\left\lVert{\texttt{sg}(z_{M+1})}\right\rVert_{2}}.$ (5)
No additional loss terms or equivariance-specific predictors are used during training.
Figure 1: seq-JEPA is a world modeling paradigm that leverages a sequence of action-observation pairs to learn architecturally distinct invariant and equivariant representations.
Figure 2: a. Transformations and observations used for training b. In predictive learning across saccades, image saliencies and inhibition of return help create more informative and less redundant patch sequences.
@@ 2.3 Action and observation sets
To evaluate generalization across transformation types, we define three sets of action-observation pairs (Figure 2). See Appendix A.1 for details of each setup.
@@
3D Invariant Equivariant Benchmark (3DIEBench). We first evaluate on the 3DIEBench dataset (Garrido et al., 2023), designed for equivariance analysis. It includes 3D object renderings with variations in rotation, floor hue, and lighting. We primarily study equivariance to $SO(3)$ rotations and secondarily to appearance factors of floor and light hue.
@@
Hand-Crafted Augmentations. In this setting, we use transformed views generated via common SSL augmentations (e.g., crop, color jitter, blur), where actions correspond to relative augmentation parameters. We use CIFAR100 and Tiny ImageNet, and follow EquiMod’s augmentation protocol (Devillers and Lefort, 2022).
Predictive Learning Across Saccades (PLS). Going beyond conventional transformations such as augmentations or 3D rotations, we show that seq-JEPA can learn visual representations from a sequence of partial observations thanks to architectural inductive biases—without relying on any hand-crafted augmentations. Our PLS has a similar flavor to I-JEPA (Assran et al., 2023) but does not require engineered masking strategies. In PLS, we train seq-JEPA on sequences of patches extracted from high-resolution images. For instance, with STL-10, we use $32\times 32$ patches to form the observation sequence. In this setting, actions correspond to the relative positions between patch centers, simulating saccadic eye movements and inducing 2-D positional equivariance to representations. To select fixation points, we adopt two biologically inspired techniques that increase informativeness, reduce redundancy, and improve the downstream utility of the aggregate representation (Figure 2):
  • Saliency-Based Fixation Sampling. Using DeepGaze IIE (Linardos et al., 2021), we extract saliency maps for each image and use them to probabilistically sample fixation points (Itti et al., 1998; Li, 2002; Zhaoping, 2014). The maps are pre-computed and introduce no training overhead.
  • Inhibition of Return (IoR). To reduce spatial overlap between patches and emulate natural exploration (Posner et al., 1985), we implement IoR by zeroing out the sampling probability of areas surrounding previously sampled fixations.